OpenStudy (anonymous):

3x+4y=7_4x+3y=10 are these lines perpendicular,parallel, or neither

OpenStudy (anonymous):

x = 19/7, y = -2/7

OpenStudy (anonymous):

neither because the slope is not the negative reciprocal both have negative slopes and the other has to be positive to be perpendicular, and they have to be the exact same to be parallel

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I probabably should have punctuated that better. Neither because the slope is not the negative reciprocal. Both have negative slopes. The equations have to be the exact opposite to be perpendicula like m=-1/2 and m=2/1 to be perpendicular, and they have to be the exact same to be parallel m=-1/2 and m=-1/2